OpenStudy (anonymous):

How would these amounts be arranged in order from greatest to least? 2/3, 4/5, 3/4, 1/8

OpenStudy (anonymous):

You need them to have a common denominator. That means that they have the same number at the bottom. Use 120

jhonyy9 (jhonyy9):

so make the first answer and after you have got the common denomintor check by with number you need multiply the numeratores after this you see what numerator divided by 120 how many will result 2/3 will be 2*40/120= 80/120 4/5 ----- 4*24/120= 96/120 3/4 ----- 3*30/120= 90/120 1/8 ----- 1*15/120= 15/120 - so and now it is very easy to arrange those all numbers in order from greatest to least